wiki:UC20180402

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 4/02/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Chip Kobulnicky (UWY), Rachel Kuzio de Naray (GSU), Mary Beth Kaiser (JHU), John Bally (CU), Joanne Hughes Clark (Seattle U), Eric Bellm (UW), Kal Kadlec (UW), Anne Verbiscer (UVa), Ben Williams (UW), Jon Holtzman (NMSU), Bill Ketzeback (APO), Conor Sayres (UW)


Identify minute-taker for this meeting: Nancy and Mark will do it

User feedback and comments from institutional representatives:

  • Wyoming - nothing to report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle U - nothing to report
  • Oklahoma - no report
  • NMSU - nothing to report
  • NAPG - no report
  • JHU - nothing to report
  • Georgia State - nothing to report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 2/28/18 through 3/29/18

0) Overview

Reasonably good weather this period but not great, lots more clouds than typical in March, we did have typical wind and dust though, a couple small storms but little moisture. DIS blue camera vacuum is being serviced this week.

1) Telescope

Telescope was operational with minimal issues. The time issues with TCC communications have been looked at and being monitored. We have updated some software and perhaps that has helped. Need more time to verify and/or investigate.

2) Instruments

DIS: Blue only improved by about 40% and quickly deteriorated so it is on the pump this week in an effort to clear any contamination. DCam continues to give us issues. Original DCam is still at Andor for servicing.

TripleSpec has been operational.

Agile has been operational this period.

Echelle is operational. There is also readout banding that is causing some issues. A replacement camera order from Fingerlakes has been ordered.

NIC-FPS is operational.

ARCTIC is operational. We are working on improving filter wheel reliability. Users should verify that their desired filter is in place. We are also seeing odd things with temperature regulation but with warming ambient temperatures things may be back to normal soon.

3) ARCSAT

M1 and M2 have been coated and should ship back within the week. Lots of testing on M2 focus seems to point to random inconsistencies. We have some suggestions for DFM on improving their software which are still being investigated.


Additional telescope and instrument discussion:

With the spring season we have experienced the usual winds over the past month. We've had a greater than normal amount of clouds, but that has not led to much precipitation. We are still way down for the year, but the National Forest Service has still been able to do some controlled burns.

Regarding the telescope and instruments, everything is working nominally.

We haven't seen the TCC issue lately (the one that was discussed last month). We are currently pumping on the DIS blue camera - that started last Friday and will pump through Thursday. Mark thinks that the contamination will not improve substantially but it will be a good data point nonetheless. John Bally asked if this implied a small leak in dewar? Mark said that the engineering staff believes there are leaks in both dewars. Generally we had been getting about a year between required pumpdowns (late spring/summer is when it gets worse), but last year and this year we are seeing it more frequently. We looked at a possible dewar redesign but this is difficult given the constraints of the DIS instrument itself (no room to spare). We also looked at getters but have not yet come up with the best plan for remediation; we expect this is not going to be an easy fix. This will likely be examined in more detail during the summer shutdown. Bally also asked if there is a helium leak detector on site, and Mark said yes, and we have used it. The results suggest a minor leak in the front o-ring but this is not a typical o-ring seal. The rest of the instruments are working pretty well.

Regarding ARCSAT, the primary and secondary are coated and on their way back to APO. In terms of the focus mechanism, Bill has been able to duplicate some of the problems we’ve seen and thinks that it is probably not mechanical in nature. It may require some software changes. Bill has looked at what it will take to collimate the mirrors once we get them back, but there is still no estimate of when ARCSAT will be back on sky.


Q2 3.5-m scheduling:

The Q2 schedule is done and there is still a little bit of OPEN time available. Russet worked hard to get everyone scheduled in terms of class trips.


Q2 ARCSAT scheduling:

There was no call issued for Q2 proposals since it is unknown when the telescope will be back up and running.


TDAWG Update

The Time Domain Astronomy Working Group (TDAWG) report is finished and is posted on the 3.5m Users' Wiki. We are now thinking about what kinds of updates we will make. One thing we’d like to do in the near term is some web page modifications to make time domain results more visible and available so that the broader ARC community is aware of what's going on and what came out of a particular target of opportunity observation.


TEWG Update

The Training and Education Working Group (TEWG) is just getting off the ground. It will be chaired by Erica Ellingson (CU), and we are still seeking representatives from several institutions (JHU and NAPG). Mary Beth said she would ask around for volunteers at JHU, and Nancy will ask for a volunteer from NAPG member institutions. The purpose of the TEWG is to examine the use of the 3.5m telescope for student training and education, specifically with the charge to: provide an overview of institutional goals and practices, identify best practices and recommendations for student training that support different institutional goals, and provide APO policy recommendations (if needed) to optimize student use. We hope to have the first TEWG telecon within the next month.


NA2 Baffle Project

John Wilson provided an email report, which Nancy relayed to the Users Committee. The current status is that the UVa team continues to compare the historical zemax files from the observatory and the unified SolidWorks models that Garrett Ebelke (UVa) has created. They also have been modifying the historical zemax model to conform to the figuring reports of the primary and secondary mirror. Currently they are considering the focal plane position in both models as its position is important since the secondary position can be moved to optimize focus at the focal plane; they expect this portion of the project to be done by next month.


KOSMOS update

There is not much to report on this front. Sarah, Kal, and Nancy have been in discussions with the KPNO staff about the compilation of the KOSMOS inventory list so that the long-term loan agreement can be completed. The KPNO staff is still working on this.


TUI Discussion

Nancy asked Conor Sayres to lead a discussion concerning TUI and the increasing prevalence of people installing it and using it on their laptops rather than using it through a single distribution that is installed on a department network. The recent problems that people have had with TUI seem to be when we have Linux users who are trying to run TUI from source. That can be problematic because TUI tries to pick up all of its dependencies from whatever packages the user has installed (e.g. numpy, matplotlib, etc.) The more recent ones have become the default for many people, and that has started to break compatibility with TUI. We don't have the manpower to make sure it is working and up to date with all versions of all packages, therefore Conor has created a standalone package for all Linux users; this is accessible from the TUI download site. Users should try the pre-packaged version of TUI first before attempting to run it from source. Jon asked shouldn't we expect TUI to work with current versions of python packages? The response was yes we should, but sometimes it can be hard to keep up. Jon suggested that the long range plan should be to migrate TUI to python 3, but in the interim the observatory needs to commit to support current versions of numpy, etc. Chip provided his perspective as his department's Linux sys admin, which is that it is fine to use a static zipped version of TUI, as long as the observatory supports its own self-contained distribution. Conor agreed with Jon, but reminded us that ultimately people want a working TUI, and miniconda does currently satisfy that (it comes with its own python and dependencies and it won’t affect users’ packages that they may have installed). Chip stated that he liked Conor’s distribution, while Jon was of the opinion that it needs to evolve (that was a problem with the old Remark software) and should be updated every 6 months or so.


ACTION ITEMS:

Action items from previous meetings:

  • Nancy will notify all 3.5m users (through an email to 35m-general) about the availability of the TDAWG report. --> DONE.
  • Nancy will begin recruiting members for the TEWG. --> DONE.

New action items from this meeting:

None.


Next meeting:

The meeting schedule for the spring semester will be the first Monday of the month at 11:00 am Mountain Time. The next call will thus be May 7, 2018.

Future meetings are scheduled for:

  • June 4
  • July 9

We will evaluate everyone's availability (as well as the telecon schedule) once we get into summer months since we will need to contend with summer travel schedules as well as summer shutdown activities.

Note: See TracWiki for help on using the wiki.